# HyperGaussians: High-Dimensional Gaussian Splatting for High-Fidelity Animatable Face Avatars

### v2

@@ 4.3 Ablation Study
As discussed in Sec. 3.2, a naive implementation of the conditioning in Eq. 4 is very inefficient for large latent codes $\boldsymbol{\gamma}_{b}$ because it requires inverting the large matrix $\boldsymbol{\Sigma}_{bb}$. The bottleneck lies in storing the covariance matrix $\boldsymbol{\Sigma}_{bb}\in\mathbb{R}^{n\times n}$. After applying the inverse covariance trick (Sec. 3.2), we only need to construct $\boldsymbol{\Sigma}_{aa}\in\mathbb{R}^{n\times n}$ by inverting $\boldsymbol{\Lambda}_{aa}^{-1}$, a small matrix. As an example, the matrix for the position attribute is only $\boldsymbol{\Lambda}_{aa}^{-1}\in\mathbb{R}^{3\times 3}$. Fig. 5 shows an empirical ablation for latent dimensionalities $n$ between $1$ and $128$. In our case study on FlashAvatar [51] (Sec. 3.3), the inverse covariance trick improves speed by $150\%$ for a small latent with $n=8$ and by $15,000\%$ for a large latent with $n=128$. Not only does the inverse covariance matrix improve speed, but it also reduces memory usage. For a small latent with $n=8$, the naïve implementation uses $42$ MB, whereas the inverse covariance trick reduces this to $22$ MB, a reduction of $48\%$. For the large latent with $n=128$, the reduction is over 90%.
Latent Dimensionality
We ablate the effect of different latent dimensionalities ($n$ in Sec. 3.2) in Tab. 3. We find that HyperGaussian are robust towards different latent dimensions. A latent dimension of $8$ performs very well, but we already observe an improvement for a single latent dimension ($n=1$) over the vanilla 3DGS variant. The bottom row corresponds to the FlashAvatar baseline [51], which doesn’t use any HyperGaussians.
Latent dim. ($n$) PSNR ↑ SSIM$\bigl{(}10^{-1}\bigr{)}$ LPIPS$\bigl{(}10^{-2}\bigr{)}$
1 29.73 9.492 5.066
2 29.92 9.503 5.000
4 29.89 9.507 4.994
8 29.99 9.510 4.978
16 29.92 9.511 4.978
32 29.89 9.511 4.976
64 29.91 9.512 5.020
Baseline 29.43 9.466 5.107
Table 3: Quantitative Ablations. Green denotes the best and Yellow the second best. The bottom row indicates the FlashAvatar [51] baseline, which does not use any HyperGaussians. We train the same model with different latent dimensions, ranging from $n=1$ to $n=64$. Employing HyperGaussians with a single latent dimension ($n=1$) already outperforms the baseline. The differences above $8$ dimensions ($n=8$) are minor, which indicates that 8 dimensions are sufficiently expressive.
@@ 5 Conclusion and Discussion
In this paper, we study how 3D Gaussian Splatting can be made more expressive for monocular face avatars. The result, HyperGaussians, is a novel extension to 3D Gaussians that offers improved expressivity and rendering quality, excelling in rendering high-frequency details like thin structures and specular reflections. Our evaluations on 19 subjects from 4 datasets outperform the state-of-the-art by simply plugging in our proposed HyperGaussians into an existing method, FlashAvatar, without any other modifications. As a limitation, HyperGaussians are high-dimensional representations and require more memory than vanilla 3DGS. Our proposed inverse covariance trick greatly reduces these requirements, but scaling HyperGaussian to thousands of dimensions is a topic for future work. In addition, HyperGaussians are not a stand-alone method, and hence, they inherit limitations of the underlying method. Future research directions could investigate the effect of HyperGaussians beyond human faces, e.g., full-body avatars or more generic dynamic scenes. In conclusion, HyperGaussians show great promise in improving high-frequency details for monocular face avatars, bringing the field a step closer to photorealistic and fast monocular face avatars.
@@ Appendix A Supplementary
This supplement contains more details and derivations in Sec. A.1, supplementary results in Sec. A.2, and discusses the potential societal impact of this work in Sec. A.3.
A.1 HyperGaussian Details
A.1.1 Formulation
This section provides more details about the formulation of HyperGaussian. We first explain how we parameterize the mean and covariance matrix. We then detail how we construct the covariance for splatting and derive the inverse covariance trick in more detail. Finally, we explain how we extract an uncertainty measure from the optimized covariance matrices.
Parameterization
@@ A.1.1 Formulation
Each HyperGaussian consists of a (typically high) dimensional mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Sigma}$ with optimizable parameters. We parameterize the mean $\boldsymbol{\mu}$ directly and decompose the covariance $\boldsymbol{\Sigma}$ into its Cholesky factor [10] $\boldsymbol{L}$, such that $\boldsymbol{\Sigma}=\boldsymbol{L}\boldsymbol{L}^{\top}$, where $\boldsymbol{L}$ is a lower triangular matrix with positive diagonal entries. To ensure the uniqueness of the factorization, we apply an exponential activation function, $\boldsymbol{L}_{i,i}(x)=e^{x}$, to the diagonal entries of the parameter matrix.
Splatting Details
This paragraph explains how we construct the covariance matrix for splatting the conditioned Gaussians [7]. It is important to note that the HyperGaussian conditional covariance $\boldsymbol{\Sigma}_{a|b}$ is not equal to the covariance $\boldsymbol{\Sigma}$ from Eq. 1. The conditional covariance $\boldsymbol{\Sigma}_{a|b}$ is used to compute the conditional mean $\boldsymbol{\mu}_{a|b}$. The covariance used for splatting is constructed from the conditional means for rotation and scale:
$\boldsymbol{\mu}_{R|\gamma_{b}}\boldsymbol{\mu}_{S|\gamma_{b}}\boldsymbol{\mu}_{S|\gamma_{b}}^{\top}\boldsymbol{\mu}_{R|\gamma_{b}}^{\top}.$ (8)
@@
The conditional covariance $\boldsymbol{\Sigma}_{a|b}$ is not actually used for splatting. We only use it for visualizing the uncertainties (see Sec. A.1.1). This is an important distinction to NDGS [10], which directly splats using the conditional matrix. NDGS has no degrees of freedom for the conditional orientation of the 3D Gaussians, i.e., the 3D Gaussians cannot rotate.
Derivation of the Inverse Covariance Trick
@@
We explain in the main paper that a naïve implementation of the conditioning is very inefficient for large latent codes $\boldsymbol{\gamma}_{b}$. Here, we provide a more detailed derivation of the covariance trick.
The bottleneck lies in storing and inverting the conditional covariance matrix $\boldsymbol{\Sigma}_{bb}\in\mathbb{R}^{n\times n}$, which tends to be a large matrix. Ideally, we would like to transfer the inversion of $\boldsymbol{\Sigma}_{bb}$ to the smaller and constant-sized block matrix $\boldsymbol{\Sigma}_{aa}$. To achieve this, we start by reformulating our HyperGaussians in terms of their precision matrix $\boldsymbol{\Lambda}=\boldsymbol{\Sigma}^{-1}$ such that $\boldsymbol{\gamma}\sim\mathcal{N}\bigl{(}\boldsymbol{\mu},\boldsymbol{\Lambda}^{-1}\bigr{)}$. We again consider the general block matrix view:
@@
This is important, as it allows us to reuse the same parameterization that we described in Sec. A.1.1 to represent $\boldsymbol{\Lambda}=\boldsymbol{L}\boldsymbol{L}^{\top}$. Conveniently, we also get the Cholesky factor $\boldsymbol{L}_{11}$ of $\boldsymbol{\Lambda}_{aa}$ as a side product:
$\begin{split}\begin{bmatrix}\boldsymbol{\Lambda}_{aa}&\boldsymbol{\Lambda}_{ab}\\ \boldsymbol{\Lambda}_{ba}&\boldsymbol{\Lambda}_{bb}\end{bmatrix}&=\begin{bmatrix}\boldsymbol{L}_{11}&\boldsymbol{0}\\ \boldsymbol{L}_{21}&\boldsymbol{L}_{22}\end{bmatrix}\begin{bmatrix}\boldsymbol{L}_{11}^{\top}&\boldsymbol{L}_{21}^{\top}\\ \boldsymbol{0}&\boldsymbol{L}_{22}^{\top}\end{bmatrix}\\ &=\begin{bmatrix}\boldsymbol{L}_{11}\boldsymbol{L}_{11}^{\top}&\boldsymbol{L}_{11}\boldsymbol{L}_{21}^{\top}\\ \boldsymbol{L}_{21}\boldsymbol{L}_{11}^{\top}&\boldsymbol{L}_{21}\boldsymbol{L}_{21}^{\top}+\boldsymbol{L}_{22}\boldsymbol{L}_{22}^{\top}\end{bmatrix}.\end{split}$ (11)
@@
Note that products with $\boldsymbol{L}_{11}^{-\top}$ can be evaluated in an efficient and numerically stable manner since $\boldsymbol{L}_{11}$ is a triangular matrix of small size, which is independent of the latent dimension. Please see the main paper for a benchmark comparison between the naïve implementation and the one applying the inverse covariance trick.
Derivation of Uncertainty
As discussed in the main paper, we observe an interesting property about HyperGaussians, which arises naturally throughout training. HyperGaussians are at their core multivariate Gaussian distributions. Their conditional covariance matrices indicate the variance of each Gaussian across the different expressions of the training subject and can be intuitively interpreted as uncertainty.

### v3

@@ 4.3 Ablation Study
We conduct ablation studies in the monocular setting (FlashAvatar [70], Sec. 3.3). We first ablate alternative means of increasing expressiveness by adding parameters to other parts of the model, and then study the computational performance gains in speed and memory when employing the inverse covariance trick (Sec. 3.2). The supp. mat. contains ablations for different latent dimensionalities.
Scaling the Model
The quality improvements provided by HyperGaussians cannot be matched by scaling other parts of the model. To demonstrate this, we increase the width and depth of the deformation MLP $F_{\Theta}$ (Fig. 2). Increasing the MLP depth deteriorates perceptual quality (LPIPS [75], lower is better) by $15\%$ (from $0.0498$ for HyperGaussians to $0.0572$ for a deeper MLP with the same number of parameters) and reduces the rendering framerate by $47\%$ from 300 FPS to 158 FPS. Increasing MLP width deteriorates perceptual quality by $3\%$ (from $0.0498$ for HyperGaussians to $0.0512$ for a wider MLP with the same number of parameters) and reduces the rendering framerate by $41\%$ from 300 FPS to 178 FPS. Please refer to the supp. mat. to see a detailed ablation for different configurations.
@@
As shown in Sec. 3.2, a naïve implementation of the conditioning in Eq. 3, as in NDGS [13], is very inefficient for large latent codes $\boldsymbol{\gamma}_{b}$, due to the large matrix $\boldsymbol{\Sigma}_{bb}\in\mathbb{R}^{n\times n}$. After applying the inverse covariance trick (Sec. 3.2), we reduce the problem to inverting the small matrix $\boldsymbol{\Lambda}_{aa}\in\mathbb{R}^{m\times m}$ ($m\in\{3,4\}$). Fig. 3 shows an empirical ablation for latent dimensionalities $n$ between 11 and $128$. In our case study on FlashAvatar [70] (Sec. 3.3), the inverse covariance trick improves speed by $150\%$ for a small latent $n=8$ and by $15,000\%$ for a large latent $n=128$. Not only does the inverse covariance matrix improve speed, but it also reduces memory usage. For a small latent $n=8$, the naïve implementation uses $42$ MB, whereas the inverse covariance trick reduces this to $22$ MB, a reduction of $48\%$. For $n=128$, it is over $90\%$.
@@ 5 Conclusion and Discussion
In this paper, we study how 3DGS can be made more expressive for monocular and multi-view face avatars. The result, HyperGaussians, is a novel extension to 3D Gaussians that offers improved expressivity and rendering quality, excelling in high-frequency details. We motivate our extension by taking a Bayesian perspective and showing that HyperGaussians perform MAP estimation over conditional distributions across 3DGS scenes. Our evaluations on 29 subjects across 6 datasets outperform the state-of-the-art by simply plugging our proposed HyperGaussians into FlashAvatar and GaussianHeadAvatar, without any other modifications. As a limitation, high-dimensional variants of 3DGS require significantly more compute and memory, which has prevented widespread adoption of NDGS. Our proposed inverse covariance trick substantially reduces these requirements and opens up future research directions that could apply HyperGaussians beyond human faces, i.e., in complex settings that require higher latent dimensions. In conclusion, HyperGaussians show great promise for improving high-frequency details, bringing the field a step closer to photorealistic and fast face avatars.
@@ Acknowledgments
This supplement contains more details and derivations in Appendices B and C, supplementary results in Appendix D, and discusses the potential societal impact of this work in Appendix E.
Appendix A Positioning with respect to Related Works
We provide an overview of the main differences between HyperGaussians and the most closely related works on Gaussian avatars in Tab. 2. We partition the methods into 3 categories: optimization-based, feed-forward, and pseudo ground-truth. Since the latter two categories require substantial computational resources for pre-training, we mainly focus on optimization-based techniques in our experiments.
Method Representation Local context Dynamic context Resource requirements
GaussianAvatars [57] 3DGS $\times$ $\times$ Low
SurFHead [31] 2DGS $\times$ $\times$ Low
SplattingAvatar [60] 3DGS $\checkmark$ $\times$ Low
MonoGaussianAvatar [10] 3DGS $\times$ $\checkmark$ Moderate
FlashAvatar (FA) [70] 3DGS $\times$ $\checkmark$ Low
GaussianHeadAvatar (GHA) [71] 3DGS $\times$ $\checkmark$ Moderate
GAGAvatar [11] FF 3DGS $\checkmark$ $\checkmark$ High
LAM [23] FF 3DGS $\checkmark$ $\times$ High
FastAvatar [39] FF 3DGS $\checkmark$ $\times$ High
Avat3r [64] FF 3DGS $\checkmark$ $\checkmark$ High
GAF [64] MVDiffusion 3DGS $\times$ $\times$ Very High
CAP4D [65] MVDiffusion 3DGS $\times$ $\times$ Very High
Ours (FA) HGS $\checkmark$ $\checkmark$ Low
Ours (GHA) HGS $\checkmark$ $\checkmark$ Moderate
Table 2: Differences to the most closely related works. GaussianAvatars [57] and SuRFHead [31] deform 3D Gaussians based on an underlying FLAME mesh [38] without local embeddings or dynamic inputs like facial expressions. SplattingAvatar [60] optimizes local embeddings, but the Gaussian properties are not dependent on expressions or pose. MonoGaussianAvatar [10], FlashAvatar [70], and GaussianHeadAvatar [71] predict expression-dependent offsets to the Gaussian properties, but their lack of local context leads to blurry or distorted results, see comparison in the main paper. Our proposed representation (HGS) attaches high-dimensional Gaussians to the mesh and optimizes learnable local embeddings for modulating the Gaussian properties based on expressions. A more recent line of work deploys large-scale generative models to either directly regress 3D Gaussian parameters or augment the training signal during optimization. GAGAvatar [11] and LAM [23] derive 3D Gaussians from DINOv2 [49] features. FastAvatar [39] and Avat3r [29] train an encoder-decoder model from scratch, where Avat3r additionally uses position maps from DUSt3R [67] and feature maps from Sapiens [27]. GAF [64] and CAP4D [65] train multi-view diffusion models [55, 61, 16] to generate pseudo ground-truth images for 3DGS optimization. These feed-forward and pseudo GT methods show promising results in sparse-view settings, their primary focus. However, they lose the ability to leverage all information when more views are available. The typical range is 1 to 4 images per forward pass. CAP4D proposes a sophisticated inference scheme to address this issue. However, it still takes 12 GPU hours to produce a single avatar. Traditional optimization-based techniques still achieve superior performance at significantly lower computational effort in such cases.
Appendix B HyperGaussian Details
This section provides more details about the formulation of HyperGaussians.
B.1 Parameterization
@@
Each HyperGaussian consists of a high-dimensional mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Sigma}$ (or precision matrix $\boldsymbol{\Lambda}$ when applying the inverse covariance trick, Sec. B.3) with optimizable parameters. We parameterize the mean $\boldsymbol{\mu}$ directly and decompose the covariance $\boldsymbol{\Sigma}$ into its Cholesky factor $\boldsymbol{L}$ [13], such that $\boldsymbol{\Sigma}=\boldsymbol{L}\boldsymbol{L}^{\top}$, where $\boldsymbol{L}$ is a lower triangular matrix with positive diagonal entries. To ensure the uniqueness of the factorization, we apply an exponential activation function, $\boldsymbol{L}_{i,i}\leftarrow\exp\boldsymbol{L}_{i,i}$, to the diagonal entries of the parameter matrix. We further show in Sec. B.3 that it is not necessary to parameterize the entire Cholesky factor $\boldsymbol{L}$ when working with the inverse covariance trick (Eqs. 11 and 12).
B.2 Splatting
This paragraph explains how we construct the 3D covariance matrix for splatting the conditioned Gaussians. We build on top of the same decomposition of $\boldsymbol{\Sigma}_{\mathrm{3D}}$ into rotation and scaling as in vanilla 3DGS, but replace the static terms $\boldsymbol{R}$ and $\boldsymbol{S}$ by their conditional analogues. More formally,
$\boldsymbol{\Sigma}_{\mathrm{3D}}=\boldsymbol{R}(\bar{\boldsymbol{q}})\bar{\boldsymbol{S}}\bar{\boldsymbol{S}}^{\top}\boldsymbol{R}(\bar{\boldsymbol{q}})^{\top},$ (7)
where $\boldsymbol{R}(\cdot)$ denotes quaternion to rotation matrix conversion, $\bar{\boldsymbol{q}}=\mathbb{E}[\mathcal{A}_{\boldsymbol{q}}|\boldsymbol{z}]$, and $\bar{\boldsymbol{S}}=\mathrm{diag}\bigl(\mathbb{E}[\mathcal{A}_{\boldsymbol{s}}|\boldsymbol{z}]\bigr)$. Moreover, $\boldsymbol{z}$ is the latent code and $\mathcal{A}$ stands for attribute (see Eq. (4) in the main paper).
@@ B.2 Splatting
This is an important distinction to NDGS [13], which directly splats using the conditional covariance matrix $\boldsymbol{\Sigma}_{\mathrm{3D}}=\mathrm{Cov}[\mathcal{A}_{\boldsymbol{\mu}_{\mathrm{3D}}}|\boldsymbol{z}]$. NDGS has no degrees of freedom for the conditional orientation of the 3D Gaussians, i.e., the 3D Gaussians cannot conditionally rotate.
To illustrate this, note that conditioning can be interpreted geometrically as taking an $m$-dimensional slice through the multivariate Gaussian with $(m+n)$ total dimensions. Fig. 7 shows two examples for $m=n=1$. Notice that the conditional covariance matrix has the same shape regardless of where the slice is taken (Eq. (3) in the main paper).
We instead use $\mathrm{Cov}[\mathcal{A}_{\boldsymbol{\mu}_{\mathrm{3D}}}|\boldsymbol{z}]$, $\mathrm{Cov}[\mathcal{A}_{\boldsymbol{q}}|\boldsymbol{z}]$, and $\mathrm{Cov}[\mathcal{A}_{\boldsymbol{s}}|\boldsymbol{z}]$ for visualizing uncertainties (see Fig. 8).
Figure 7: Geometric interpretation of Gaussian conditioning on two examples with large (left) and small (right) uncertainty at different realizations of $\boldsymbol{\gamma}_{b}$. As a result, the conditional mean shifts, while the conditional covariance is the same for both slices.
B.3 Derivation of the Inverse Covariance Trick
@@
We explain in the main paper that a naïve implementation of the conditioning is very inefficient for large latent codes $\boldsymbol{\gamma}_{b}$. Here, we provide a more detailed derivation of the inverse covariance trick.
We reformulate our HyperGaussians in terms of their precision matrix $\boldsymbol{\Lambda}=\boldsymbol{\Sigma}^{-1}$ such that $\boldsymbol{\gamma}\sim\mathcal{N}\bigl(\boldsymbol{\mu},\boldsymbol{\Lambda}^{-1}\bigr)$. We consider the following block matrix view:
@@
This is important, as it allows us to reuse the same parameterization that we described in Sec. B.1 to represent $\boldsymbol{\Lambda}=\boldsymbol{L}\boldsymbol{L}^{\top}$. Conveniently, we also get the Cholesky factor $\boldsymbol{L}_{11}$ of $\boldsymbol{\Lambda}_{aa}$ as a side product:
[…]
@@
Note that products with $\boldsymbol{L}_{11}^{-\top}$ can be evaluated in an efficient and numerically stable manner since $\boldsymbol{L}_{11}$ is a triangular matrix of small size, which is independent of the latent dimension. These equations also show that it is sufficient only to parameterize $\boldsymbol{L}_{11}$ and $\boldsymbol{L}_{21}$ for modeling the conditional distribution. This reduces the number of parameters from $\mathcal{O}\bigl((m+n)^{2}\bigr)$ to $\mathcal{O}\bigl(m^{2}+mn\bigr)$ compared to the full parameterization. Please see the main paper for a benchmark comparison between the naïve implementation and the one applying the inverse covariance trick.
B.4 Derivation of Uncertainty
We observe an interesting property about HyperGaussians, which arises naturally from their Bayesian interpretation. HyperGaussians are at their core multivariate Gaussian distributions. Their conditional covariance matrices indicate the variance of each Gaussian across the different expressions of the training subject and can be intuitively interpreted as uncertainty.
@@ B.4 Derivation of Uncertainty
$\begin{split}\begin{aligned} \sigma:=&\log\det\boldsymbol{\Sigma}_{a|b}\\ =&-\log\det\boldsymbol{\Lambda}_{aa}&&(1)\\ =&-2\log\det\boldsymbol{L}_{11}&&(2)\\ =&-2\operatorname{tr}\log\boldsymbol{L}_{11},&&(3)\end{aligned}\end{split}$ (13)
where we used $\det\boldsymbol{\Sigma}_{a|b}=\det\boldsymbol{\Lambda}_{aa}^{-1}=\bigl(\det\boldsymbol{\Lambda}_{aa}\bigr)^{-1}$ in step $(1)$, $\det\boldsymbol{\Lambda}_{aa}=\det\boldsymbol{L}_{11}\boldsymbol{L}_{11}^{\top}=\bigl(\det\boldsymbol{L}_{11}\bigr)^{2}$ in step $(2)$, and $\det\boldsymbol{L}_{11}=\prod_{i=1}^{m}(\boldsymbol{L}_{11})_{i,i}$ in step $(3)$. The $\log$ in step $(3)$ is applied element-wise.
Again, the inverse covariance trick (Sec. B.3) is key to efficiently compute this quantity. These values are summed up across the conditional distributions for all Gaussian attributes. To render these uncertainties, we further apply a sigmoid function and map the values to colors. This agreement between the uncertainty estimates and what would intuitively be considered difficult regions emerges without explicit supervision. We demonstrate an example of this effect in Fig. 8. Note that the foreground mask in the bottom-left neck area was unstable throughout the video, leading to high uncertainty despite being rigid. Moreover, the uncertainty correctly captures the variability of the lips and specular reflections on glasses. We further observe that, while rigid, the glass frames exhibit significant displacements due to underlying mesh deformations.
Figure 8: Uncertainty quantification on one of the training subjects. Green denotes low uncertainty, while Red denotes high uncertainty. Note that the semantic structure arises purely from the probabilistic formulation without additional supervision.
Appendix C GaussianHeadAvatar Details
GaussianHeadAvatar (GHA) [71] represents a scene with $N$ Gaussians, which have position $X$, multi-channel color $C$, rotation $Q$, scale $S$ and opacity $A$. The way GHA models expression-dependent effects is by maintaining a canonical set of Gaussians $\{X_{0},F_{0},Q_{0},S_{0},A_{0}\}$, where $F_{0}$ is a per-point feature vector, and training an MLP-based expression conditioning dynamic generator $\Phi$, which predicts dynamic changes with respect to the canonical model. More specifically, given expression $\theta$ and head pose $\beta$, they compute
$\{X,C,Q,S,A\}=\Phi(X_{0},F_{0},Q_{0},S_{0},A_{0};\theta,\beta)$ (14)
in order to obtain a deformed set of Gaussians.
For our HyperGaussians integration, we modify the pipeline such that the MLP outputs per-Gaussian latent codes for each attribute, which are then used to condition the HyperGaussians. Formally,
$\{\boldsymbol{z}_{X},\boldsymbol{z}_{C},\boldsymbol{z}_{Q},\boldsymbol{z}_{S},\boldsymbol{z}_{A}\}=\Phi(X_{0},F_{0},Q_{0},S_{0},A_{0};\theta,\beta)$ (15)
are the expression-dependent latents, which are then used for computing the conditional means $\mathbb{E}[\mathcal{A}|\boldsymbol{z}]$ that are fed to the remainder of the pipeline, just like the MLP offsets in the original method, analogous to our FlashAvatar integration.
Appendix D Supplementary Experiments
NeRSemble Details
For our multi-view setting, we used 10 subjects from the NeRSemble [28] dataset, with the following IDs: 017, 018, 024, 031, 033, 036, 037, 129, 141, 144. We select all sequences labelled with EMO, EXP (excluding tongue sequences), SEN for training, and the FREE sequences for testing.
D.1 Qualitative Results
We show video results for self- and cross-reenactment on the supplementary HTML page. In addition, Figs. 9 and 10 show additional results against NDGS [13] after integration into the baseline methods, similar to HyperGaussians. Moreover, Fig. 11 provides qualitative results for varying latent dimensionalities.
Figure 9: Comparison with NDGS integrated into FlashAvatar. The limited degrees of freedom in NDGS lead to misalignments of thin structures and edges. The numbers show PSNR, SSIM ($10^{-1}$), and LPIPS ($10^{-2}$).
D.2 Comparison with NDGS
Our approach differs from NDGS [13] in its mathematical formulation and capabilities. While NDGS uses the conditional covariance matrix, which is independent of the latent code, to represent the size and shape directly, our HyperGaussians apply multivariate Gaussians on each attribute where the conditional means dynamically adapt the location, scale, and orientation of the derived 3D Gaussians in response to the latent code. This crucial difference gives HyperGaussians the necessary degrees of freedom to model thin geometry and complex deformations with higher accuracy. Fig. 9 shows our method eliminates the artifacts visible in NDGS (particularly on glass frames) and produces substantially more precise geometry at curved boundaries. We further demonstrate the benefits of HyperGaussians over NDSG by integrating them into GaussianHeadAvatar [71] in a similar fashion. Again, we can observe in Fig. 10 that NDGS produces inferior reconstructions compared to HyperGaussians. More specifically, it lacks detail for specular reflections on the glasses and teeth (row 3), and it fails to accurately position the glasses at the correct vertical location (a downward shift in row 4 due to surrounding skin deformations). Moreover, the bottom-right wrinkles in row 4 disappear when the expression latent deviates too far from the mean, since NDGS uses the joint probability density to modulate opacity. These observations can also be made quantitatively in Tab. 3.